Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night Report 2016-03-07

Observing: Stefan Taubenberger, Tamar Faran, Markus Kromer

Support Team: N. Elias-Rosa

TAT: Joe Anderson, Annalisa De Cia


Weather conditions


Start of night: thick clouds, humidity ~ 85%, wind~ 5 m/s

Middle of night: clear, humidity ~ 65%, wind ~10 m/s

End of night: clear, humidity ~ 50%, wind ~12 m/s


------------------------------------------------

Calibrations


Evening: PESSTO_SOFI_SpecArc, PESSTO_SOFI_specflat *

Morning:


*JHK dome flats as in PESSTO_SOFI_FlatJHK were taken by the TO under a different program ID and can be found in 59.A-9002(D)


------------------------------------------------


** Problems in the beginning of the night

- Clouds and humidity in the beginning of the night prevented the start of the observations.

- A problem occurred with the shutter of EFOSC, which could not be fixed during the night. At 03:00 UT switched to SOFI and started observations at 05:00 UT after the humidity level went down.


05:10 UT: Follow up: PS15br, JHK, 1080s + 1440s + 1440s, seeing ~1.4"

06:40 UT: Follow up, SN2016aiy, 400s + 360s + 720s, JHK, seeing ~1"

07:15 UT: Follow up: S2016adj, BG + RG, 18m, seeing ~1" *

08:10 UT: Telluric Standard, HIP64550, BG + RG, 7s, seeing ~1"

08:15 UT: Follow up: LSQ15adm, JHK, 1080s + 1440s + 1440s, seeing ~1.6"

09:45 UT: Follow up: SN2016adj, JHK, 10m, seeing ~1"

09:50 UT: Standard: EG274, BG + RG, ~10m, seeing ~0.9"


10:00 UT: End of the night


* The exposure took longer than originally calculated, due to a problem possibly related to the readout time of SOFI. Therefore, observation of the RG had to be stopped in the middle in order to complete the rest of the requests.



Comments